(3 points) Complete the table below Books read within the past year Frequency Relative Frequency Cumulative Relative Frequency none 1-5 6- 10 11-15 16- 20 21 -25 Total 0.166666666666667 0.2 її
Answer #1

0.277777... = 5/18 = 10/36

0.16666.. = 1/6 = 6/36

total can be 36

by trial and error

f relative cumulative
none 1 0.027778 0.027778
1,5 6 0.166667 0.194444
6,10 1 0.027778 0.222222
11,15 10 0.277778 0.5
16,20 11 0.305556 0.805556
21,25 7 0.194444 1
total 36 1